OpenStudy (anonymous):

)A bag contains 40 marbles, 4 of which are blue, 10 are red, 25 are green, and 1 is purple. Shawna takes a marble out of the bag, records the color, and returns it to the bag. How many green marbles should she expect after 400 trials? 25 40 100 250

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (anonymous):

i will fan and medal

OpenStudy (pinklion23):

maybe i can help

OpenStudy (anonymous):

ok

OpenStudy (bloomlocke367):

okay... you need to find out how many green marbles there are and how many total marbles there are...

OpenStudy (anonymous):

there is 25 green and 40 total

OpenStudy (bloomlocke367):

well it's already there, but you get the point XD

OpenStudy (bloomlocke367):

you could write a proportion...

OpenStudy (anonymous):

there is 25 over 40 right

OpenStudy (anonymous):

25 out of 40

OpenStudy (bloomlocke367):

\[\frac{ 25 }{ 40 }=\frac{? }{400}\] yep!

OpenStudy (bloomlocke367):

do you know how to solve for the '?'

OpenStudy (anonymous):

i multiply right

OpenStudy (bloomlocke367):

yes.. you cross-multiply and divide :)

OpenStudy (anonymous):

thanks

OpenStudy (bloomlocke367):

what do you get?

OpenStudy (anonymous):

have a good day bloom

OpenStudy (bloomlocke367):

you too. \(\Huge\mathbb{\color{lime}{GREAT~WORK}}\)

OpenStudy (anonymous):

thanks bye